In a certain class, there are 14 boys and 21 girls. If the average score on an exam is 92 and the boys' average score is 89, what is the girls' average score?
Answer by alicealc(293)

(14 * 89 + 21 * x )/(14 + 21) = 92
(1246 + 21x)/35 = 92
1246 + 21x = 92 * 35
1246 + 21x = 3220
21x = 3220 - 1246
21x = 1974
x = 1974/21 = 94

so, the girls' average score is 94